# Robust Control Barrier Functions using Uncertainty Estimation with Application to Mobile Robots

### v2

@@ III-C Uncertainty Estimation-Based Safety with HOCBFs
$\begin{split}&\!\!\tilde{a}(t,\!x)\triangleq L_{\hat{g}}\Psi(x)\!+\!\big{(}\hat{g}^{\top}\!(x)\mathcal{F}(x)\hat{\Delta}(t)\big{)}^{\top},\\ &\!\!\tilde{b}(t,\!x)\triangleq L_{\hat{f}}\Psi(x)\!+\!L_{\hat{g}}\Psi(x)\big{(}\!-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}\!+\!\nabla\Psi(x)\hat{\Delta}(t)\\ &\!\!+\big{(}\hat{f}(x)+\hat{g}(x)\big{(}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}+\hat{\Delta}(t)\big{)}^{\top}\mathcal{F}(x)\hat{\Delta}(t)\\ &\!\!-\big{\|}\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\big{\|}\delta_{L}-\Omega(t,x,e)\\ &\!\!+\alpha_{r}\!\big{(}\Psi(x)\!+\!\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\hat{\Delta}(t)\!-\!\big{\|}\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\big{\|}\bar{e}(t)\big{)}\\ &\!\!\tilde{c}(t,\!x,\!u)\!\triangleq\!\bar{e}(t)\big{\|}\big{(}\hat{f}(x)\!+\!\hat{g}(x)\big{(}\bar{u}\!-\!\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}\!+\!\hat{\Delta}(t)\big{)}^{\top}\!\mathcal{F}(x)\big{\|}.\end{split}$ (37)
@@
$\begin{array}[]{llll}{\mathbf{\bar{k}^{*}}(t,x)=\ }\displaystyle\operatorname*{arg\,min}_{\bar{u}\in\mathbb{R}^{m},~{}\rho\in\mathbb{R}}\ \ \ \ {\rho}\\[11.38109pt] ~{}~{}~{}~{}\textrm{s.t.}\ \ \tilde{c}(t,x,u)\leq\tilde{a}(t,x)\bar{u}+\tilde{b}(t,x)\\[2.84526pt] ~{}~{}~{}~{}~{}~{}~{}~{}~{}\|\bar{u}-\mathbf{k_{d}}(x)\|^{2}\leq\rho\\[2.84526pt] ~{}~{}~{}~{}~{}~{}~{}~{}~{}\bar{u}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\in U,\end{array}$
@@
The UE-HOCBF-SOCP is the solution to optimization problem above with the second constraint replaced by (38). The UE-HOCBF-SOCP guarantees forward invariance of ${\mathcal{C}(t)}$.
@@
Specifically, if ${(\hat{g}(x)\bar{u})^{\top}\!\mathcal{F}(x)\!\equiv\!0}$, the left-hand side of the inequality in the UE-HOCBF safety filter becomes independent of $\bar{u}$, and the optimization problem reduces to a QP. More specifically, suppose that ${\mathcal{F}(x)\!\equiv\!0}$, then the constraint can be integrated into an uncertainty estimation-based QP, UE-HOCBF-QP. Furthermore, if ${(\hat{g}(x)\bar{u})^{\top}\!\mathcal{F}(x)\!\equiv\!0}$, we can extend Theorem 4 to HOCBFs. In order to achieve this objective, we need to modify ${\phi_{r-1}}$ and its 0-superlevel set to provide robustness against estimation error $e$, analogous to (31):
@@
IV Simulations And Hardware Experiments
IV-A Elastic Actuator with Matched and Unmatched Uncertainties
@@
We choose a CBF, ${h(x)\!=\!x_{2}\!-\!c_{q}x_{3}}$, where ${c_{q}\!=\!-2}$, and ${\alpha_{1}(h)\!=\!2h,\alpha_{2}(h)\!=\!2h}$. The CBF ${h}$ has IRD of 2, i.e., ${r\!=\!2}$ in (II-B), and DRD of 1. This uncertainty structure and CBF choice satisfy Assumption 3. We use an HOCBF framework. From (LABEL:eq:secphi) and (III-C) we have ${\phi_{0}(x)\!=\!x_{2}\!-\!c_{q}x_{3}}$, ${\phi_{1}(x)\!\!=\!\!\Delta_{2}(x)\!+\!{{m\bar{g}L\sin{{x}_{1}}\!-\!k({x}_{1}\!-\!{x}_{3})}}/{I_{L}}\!-\!c_{q}x_{4}\!+\!\alpha_{1}(x_{2}\!-\!c_{q}x_{3}\!)}$, and ${\mathcal{F}(x)\!\equiv\!0}$. Since ${\mathcal{F}(x)\!\equiv\!0}$, we use an UE-HOCBF-QP as a safety filter. Substituting ${\phi_{0},~{}\phi_{1}}$ into (LABEL:eq:qcqpparam) yields an affine HOCBF condition for the actuator system for some ${\alpha_{2}\!\in\!\mathcal{K}_{\infty,e}}$. The initial state values are ${x(0)\!=\![0~{}\!0.5~{}\!0~{}\!\!-\!0.2]^{\top}}$. To obtain bounding constants ${\delta_{L},\delta_{b}}$ for the problem’s uncertainty, we simulate numerous scenarios, resulting in ${\delta_{L}\!=\!0.6,\delta_{b}\!=\!1.4}$. We choose ${\Lambda\!=\!\operatorname{diag}(5,5,5,5)}$ resulting in ${\mu_{e}\!=\!1.25}$, and ${H\!=\!\operatorname{diag}(1,1,1,1)}$ to solve the Lyapunov equation.
Figure 2: Simulations of the uncertain elastic actuator. (Left) (Top) State and reference tracking of the uncertain system operating under the UE-HOCBF-QP. (Bottom) Values of the safe control input, $u$ safe, and the nominal input, $u$ nominal, vs. time. (Center) (Top) The proposed estimator can effectively estimate actual model uncertainties within the quantified bounds. Because of ${\hat{\Delta}(0)\!=\!\bf 0}$ and ${\Delta_{1}\!\equiv\!0,~{}\Delta_{3}\!\equiv\!0}$, the estimator output is zero for these uncertainty components. (Center) (Bottom) The estimation error with the theoretical bound which is satisfied. (Right) (Top) The value of CBF $h$ vs. time and trajectory of the system with the boundary of the safe set (Bottom). The proposed UE-HOCBF-QP safety filter maintains safety in the presence of unmodeled dynamics. While HOCBF-QP satisfies safety for the nominal system, it leaves the safe set with the uncertain system. Under the robustified HOCBF (R-HOCBF) approach [30], the system leaves the safe set due to ignoring the unmatched uncertainty, which leads to safety violations.
We compared the proposed method to the robustified HOCBF [22] (Remark 3) with a known time-invariant upper bound of the matched uncertainty ${\Delta_{m}}$ such that ${\|\Delta_{m}(x,u)\|\!<\!\bar{\Delta}_{m}}$, ${\forall x\!\in\!X}$, ${\forall u\!\in\!U}$. However, this method ignores the unmatched uncertainty ${\Delta}_{u}$ and thus does not guarantee safety in the presence of unmatched disturbances, as it relies on the assumption that ${\text{IRD}\!=\!\text{DRD}\!=\!r}$. As discussed in Section I, our method is, to the best of our knowledge, the first to ensure robust safety against both matched and unmatched uncertainties via HOCBFs.
@@
IV-B Experimental Validation of Mobile Robot Example
@@
Figure 3: Photo of a tracked mobile robot on an inclined surface during one experiment. The higher friction carpet patches on the slope result in slip uncertainty when only one track lies on a patch (Left). The geometry of full-state matched and unmatched uncertainties due to skidding and slipping, and description of the safe set ${\mathcal{C}}$ for Example IV-B (Right).
@@
Next, we consider how the uncertainty $\Delta$ in (39) impacts the CBF constraint. The mobile robot must maintain a safe distance ${\tau\!\in\!\mathbb{R}^{+}_{0}}$ from a wall aligned in the ${y}^{I}$ direction: $h(x)\!=\!{y}^{I}\!+\!\xi_{\theta}\sin{\theta}\!+\!\tau$, where ${\xi_{\theta}\!\in\!\mathbb{R}^{+}_{0}}$ determines the effect of the heading angle on safety, (see Fig. 3). This CBF choice results in $L_{\hat{f}}h(x)\!=\!0$, $L_{\hat{g}}h(x)\!=\![\sin{\theta}~{}\xi_{\theta}\cos{\theta}]^{\top}$, $\nabla h\Delta(t\!,x,\!u)\!=\!v(\sin{\gamma}\!-\!\sin{\theta})\!-\!d_{x_{B}}\!\sin{\theta}\!+\!d_{\theta}\xi_{\theta}\cos{\theta}$, and the CBF ${h}$ has ${\text{DRD}\!=\!\text{IRD}\!=\!1}$.
@@
The left-hand side of (LABEL:eq:cbfununi) is satisfied since $h$ is a CBF for the nominal system. The right-hand side of (LABEL:eq:cbfununi) quantifies the effect of uncertainty on safety: when it is positive (e.g., if ${v\!>\!0}$, ${-\pi/4\!<\!\theta\!<\!0}$, ${-\pi/4\!<\!\beta\!<\!0}$, ${d_{x_{B}}\!\leq\!0}$, and ${d_{\theta}\!\leq\!0}$, then the right-hand term will be positive), safety may be violated. To address this robustness problem, we implemented UE-CBF-QP.
@@
The reference trajectory of the robot is chosen to violate the safety when using the nominal controller. Under a CBF-QP approach, the robot’s path leaves the safe set due to the unmodeled uncertainty–see Fig. 4 (Right) (Middle) and (Bottom). Note that the uncertainty vector is the difference between the discrete-time derivative of the hardware system’s flow map and the nominal robot model. Under the proposed method, the robot states remain within the 0-superlevel set of the chosen CBF. Fig. 4 (Center) shows that our approach can estimate the model uncertainty, which affects the safety constraints.
Figure 4: Experimental results for the mobile robot example of Fig. 3. (Left) (Top) State trajectories (${x^{I}~{}[m],~{}y^{I}~{}[m],~{}\theta~{}[rad]}$) and goal position of a robot operating under the UE-CBF-QP. (Middle) and (Bottom) The values of the safe linear velocity input, $v~{}[m/s]$ safe, the angular velocity input, $\omega~{}[rad/s]$ safe, and the nominal control inputs, $v$ nominal, $\omega$ nominal, vs. time. (Center) The uncertainty estimator tracks modeling inaccuracies, as shown by the first and last components of the uncertainty vector, ${\Delta_{1}}$ and ${\Delta_{3}}$. (Right) (Top) The uncertainty estimation error stays within the theoretical bounds. (Middle) The value of CBF $h$ vs. time (Bottom) System trajectory is plotted with the safe set boundary. The proposed UE-CBF-QP-based safety filter maintains a safe distance from an edge on a slope in the presence of unmodeled dynamics. However, the CBF-QP controlled robot leaves the safe set due to the slope-induced uncertainty.
V Conclusions and Future Work
We presented a method to synthesize robust controllers with formal safety guarantees for nonlinear systems that experience both matched unmatched uncertainty. A novel uncertainty estimator is applied to state and input-dependent uncertainty, leading to bound in the CBF framework. By integrating the estimator with CBFs and HOCBFs, robust safety conditions were derived. The utility of our frameworks was experimentally demonstrated by controlling a tracked mobile robot on a slope while using a reduced-order nominal model.
Future work will eliminate the choice of the bound $\delta_{L}$ via data-driven techniques. Additionally, our methods can also consider state estimation errors.

### v3

@@ 3.3 Uncertainty Estimation-Based Safety with HOCBFs
$\begin{split}&\!\!\tilde{a}(t,\!x)\triangleq L_{\hat{g}}\Psi(x)\!+\!\big{(}\hat{g}^{\top}\!(x)\mathcal{F}(x)\hat{\Delta}(t)\big{)}^{\top},\\ &\!\!\tilde{b}(t,\!x)\triangleq L_{\hat{f}}\Psi(x)\!+\!L_{\hat{g}}\Psi(x)\big{(}\!-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}\!+\!\nabla\Psi(x)\hat{\Delta}(t)\\ &\!\!+\big{(}\hat{f}(x)+\hat{g}(x)\big{(}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}+\hat{\Delta}(t)\big{)}^{\top}\mathcal{F}(x)\hat{\Delta}(t)\\ &\!\!-\big{\|}\nabla L_{\hat{f}}^{r-2}h(x)\big{\|}\delta_{L}-\Omega(t,x,e)\\ &\!\!+\alpha_{r}\!\big{(}\Psi(x)\!+\!\nabla L_{\hat{f}}^{r-2}h(x)\hat{\Delta}(t)\!-\!\big{\|}\nabla L_{\hat{f}}^{r-2}h(x)\big{\|}\bar{e}(t)\big{)}\\ &\!\!\tilde{c}(t,\!x,\!u)\!\triangleq\!\bar{e}(t)\big{\|}\big{(}\hat{f}(x)\!+\!\hat{g}(x)\big{(}\bar{u}\!-\!\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}\!+\!\hat{\Delta}(t)\big{)}^{\top}\!\mathcal{F}(x)\big{\|}.\end{split}$ (37)
@@
$\begin{array}[]{llll}{[\mathbf{\bar{k}^{*}}(t,x),~{}\rho^{*}]=\ }\displaystyle\operatorname*{arg\,min}_{\bar{u}\in\mathbb{R}^{m},~{}\rho\in\mathbb{R}}\ \ \ \ {\rho}\\[11.38109pt] ~{}~{}~{}~{}\textrm{s.t.}\ \ \tilde{c}(t,x,u)\leq\tilde{a}(t,x)\bar{u}+\tilde{b}(t,x)\\[2.84526pt] ~{}~{}~{}~{}~{}~{}~{}~{}~{}\|\bar{u}-\mathbf{k_{d}}(x)\|^{2}\leq\rho\\[2.84526pt] ~{}~{}~{}~{}~{}~{}~{}~{}~{}\bar{u}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\in U,\end{array}$
@@
The UE-HOCBF-SOCP is the solution to the optimization problem above with the second constraint replaced by (38). The UE-HOCBF-SOCP guarantees forward invariance of ${\mathcal{C}(t)}$.
@@
If ${(\hat{g}(x)\bar{u})^{\top}\!\mathcal{F}(x)\!\equiv\!0}$, the left-hand side of the inequality in the UE-HOCBF safety filter becomes independent of $\bar{u}$, and the optimization problem reduces to a QP. More specifically, suppose that ${\mathcal{F}(x)\!\equiv\!0}$, then the constraint can be integrated into an uncertainty estimation-based QP, UE-HOCBF-QP. Furthermore, if ${(\hat{g}(x)\bar{u})^{\top}\!\mathcal{F}(x)\!\equiv\!0}$, we can extend Theorem 4 to HOCBFs. In order to achieve this objective, we need to modify ${\phi_{r-1}}$ and its 0-superlevel set to provide robustness against estimation error $e$, analogous to (31):
@@
4 Simulations And Hardware Experiments
4.1 Elastic Actuator with Matched and Unmatched Uncertainties
@@
We choose a CBF, ${h(x)\!=\!x_{2}\!-\!c_{q}x_{3}}$, where ${c_{q}\!=\!-2}$, and ${\alpha_{1}(h)\!=\!2h,\alpha_{2}(h)\!=\!2h}$. The CBF ${h}$ has IRD of 2, i.e., ${r\!=\!2}$ in (2.2), and DRD of 1. This uncertainty structure and CBF choice satisfy Assumption 3. We use an HOCBF framework. From (LABEL:eq:secphi) and (3.3) we have ${\phi_{0}(x)\!=\!x_{2}\!-\!c_{q}x_{3}}$, ${\phi_{1}(x)\!\!=\!\!\Delta_{2}(x)\!+\!{{m\bar{g}L\sin{{x}_{1}}\!-\!k({x}_{1}\!-\!{x}_{3})}}/{I_{L}}\!-\!c_{q}x_{4}\!+\!\alpha_{1}(x_{2}\!-\!c_{q}x_{3}\!)}$, and ${\mathcal{F}(x)\!\equiv\!0}$. Since ${\mathcal{F}(x)\!\equiv\!0}$, we use an UE-HOCBF-QP as a safety filter. Substituting ${\phi_{0},~{}\phi_{1}}$ into (LABEL:eq:qcqpparam) yields an affine HOCBF condition for the actuator system for some ${\alpha_{2}\!\in\!\mathcal{K}_{\infty,e}}$. The initial state values are ${x(0)\!=\![0~{}\!0.5~{}\!0~{}\!\!-\!0.2]^{\top}}$. To obtain bounding constants ${\delta_{L},\delta_{b}}$ for the problem’s uncertainty, we simulate numerous scenarios, resulting in ${\delta_{L}\!=\!0.6,\delta_{b}\!=\!1.4}$. We choose ${\Lambda\!=\!\operatorname{diag}(5,5,5,5)}$ resulting in ${\mu_{e}\!=\!1.25}$, and ${H\!=\!\operatorname{diag}(1,1,1,1)}$ to solve the Lyapunov equation.
Figure 2: Simulations of the uncertain elastic actuator. (Left) (Top) State and reference tracking of the uncertain system operating under the UE-HOCBF-QP. (Bottom) Values of the safe control input, $u$ safe, and the nominal input, $u$ nominal, vs. time. (Center) (Top) The proposed estimator can effectively estimate actual model uncertainties within the quantified bounds. Because of ${\hat{\Delta}(0)\!=\!\bf 0}$ and ${\Delta_{1}\!\equiv\!0,~{}\Delta_{3}\!\equiv\!0}$, the estimator output is zero for these uncertainty components. (Bottom) The estimation error with the theoretical bound which is satisfied. (Right) (Top) The value of CBF $h$ vs. time and trajectory of the system with the boundary of the safe set. (Bottom) The proposed UE-HOCBF-QP safety filter maintains safety in the presence of unmodeled dynamics. While HOCBF-QP satisfies safety for the nominal system, it leaves the safe set with the uncertain system. Under the robustified HOCBF (R-HOCBF) approach [22], the system leaves the safe set due to ignoring the unmatched uncertainty, which leads to safety violations.
We compared the proposed method to the robustified HOCBF [22] (Remark 3) with a known time-invariant upper bound of the matched uncertainty ${\Delta_{m}}$ such that ${\|\Delta_{m}(x,u)\|\!<\!\bar{\Delta}_{m}}$, ${\forall(x,u)\in X\times U}$. However, this method ignores the unmatched uncertainty ${\Delta}_{u}$ and thus does not guarantee safety in the presence of unmatched disturbances, as it relies on the assumption that ${\text{IRD}\!=\!\text{DRD}\!=\!r}$. As discussed in Section 1, our method is, to the best of our knowledge, the first to ensure robust safety against both matched and unmatched uncertainties via HOCBFs.
@@
4.2 Experimental Validation on a Mobile Robot
@@
Figure 3: Photo of a tracked mobile robot on an inclined surface during one experiment. The higher friction carpet patches on the slope result in slip uncertainty when only one track lies on a patch (Left). The geometry of full-state matched and unmatched uncertainties due to skidding and slipping, and description of the safe set ${\mathcal{C}}$ for Example 4.2 (Right).
@@
Next, we consider how the uncertainty $\Delta$ in (39) impacts the CBF constraint. Assume that the mobile robot must maintain a safe distance ${\tau\!\in\!\mathbb{R}^{+}_{0}}$ froma wall aligned in the ${y}^{I}$ direction: $h(x)\!=\!{y}^{I}\!+\!\xi_{\theta}\sin{\theta}\!+\!\tau$, where ${\xi_{\theta}\!\in\!\mathbb{R}^{+}_{0}}$determines the effect of the heading angle on safety (see Fig. 3). This CBF choice results in $L_{\hat{f}}h(x)\!=\!0$, $L_{\hat{g}}h(x)\!=\![\sin{\theta}~{}\xi_{\theta}\cos{\theta}]^{\top}$, $\nabla h\Delta(t\!,x,\!u)\!=\!v(\sin{\gamma}\!-\!\sin{\theta})\!-\!d_{x_{B}}\!\sin{\theta}\!+\!d_{\theta}\xi_{\theta}\cos{\theta}$, and the CBF ${h}$ has ${\text{DRD}\!=\!\text{IRD}\!=\!1}$.
@@
The left-hand side of (LABEL:eq:cbfununi) is satisfied since $h$ is a CBF for the nominal system. The right-hand side of (LABEL:eq:cbfununi) quantifies the effect of uncertainty on safety: when it is positive (e.g., if ${v\!>\!0}$, ${-\pi/4\!<\!\theta\!<\!0}$, ${-\pi/4\!<\!\beta\!<\!0}$, ${d_{x_{B}}\!\leq\!0}$, and ${d_{\theta}\!\leq\!0}$, then the right-hand term will be positive), safety may be violated. To address this robustness problem, we implemented our UE-CBF-QP.
@@
The robot’s reference trajectory is chosen to violate the safety constraint when using the nominal controller. Under a CBF-QP approach, the robot’s path leaves the safe set due to the unmodeled uncertainty–see Fig. 4 (Right) (Middle) and (Bottom). Note that the uncertainty vector is the difference between the discrete-time derivative of the hardware system’s flow mapand the nominal robot model. Uthe robot states remainwithin the 0-superlevel set of the chosen CBF. Fig. 4 (Center) shows that our approach can estimate the model uncertainty, which affects the safety constraints.
Figure 4: Experimental results for the mobile robot example of Fig. 3. (Left) (Top) State trajectories (${x^{I}~{}[m],~{}y^{I}~{}[m],~{}\theta~{}[rad]}$) and goal position of a robot operating under the UE-CBF-QP. (Middle) and (Bottom) The values of the safe linear velocity input, $v~{}[m/s]$ safe, the angular velocity input, $\omega~{}[rad/s]$ safe, and the nominal control inputs, $v$ nominal, $\omega$ nominal, vs. time. (Center) The uncertainty estimator tracksmodeing inaccuracies, as shown by the first and last components of the uncertainty vector,${\Delta_{1}}$ and ${\Delta_{3}}$. (Right) (Top) The uncertainty estimation error stays within the theoretical bounds. (Middle) The value of CBF $h$ vs. time. (Bottom) System trajectory is plotted with the safe set boundary. The proposed UE-CBF-QP-based safety filter maintains a safe distance from an edge on a slope in the presence of unmodeed dynamics. However, the CBF-QP controlled robot leaves the safe set due to the slope-induced uncertainty.
5 Conclusions and Future Work
We presented a method to synthesize robust controllers with formal safety guarantees for nonlinear systems that experience both matched and unmatched uncertainty. A novel uncertainty estimator is applied to state and input-dependent uncertainty, leading to bound in the CBF framework. By integrating the estimator with CBFs and HOCBFs, robust safety conditions were derived. The utility of ourframeworks was experimentally demonstrated by controlling a tracked mobile robot on a slope while using a reduced-order nominal model.
Future work will estimate the bound $\delta_{L}$ via data-driven techniques. Additionally, our methods can also consider state estimation errors.